OpenStudy (anonymous):

If f(-2) = 10, then f(x)= 3x + ____?

jimthompson5910 (jim_thompson5910):

I'm going to use k in place of the blank spot f(x)= 3x + k f(-2)= 3(-2) + k ... plug in x = -2 (ie replace every copy of x with -2) 10= 3(-2) + k ... plug in f(-2) = 10 (ie replace every copy of f(-2) with 10) Now we solve for k 10= 3(-2) + k 10= -6 + k 10+6 = k 16 = k k = 16 which means that f(x)= 3x + k turns into f(x)= 3x + 16

OpenStudy (anonymous):

thank you, you really made me understand.

jimthompson5910 (jim_thompson5910):

You're welcome. I'm glad to be of help.

OpenStudy (anonymous):

one question, what did you add to get 10 ?

jimthompson5910 (jim_thompson5910):

the 10 is given in the portion where it says "f(-2) = 10"

jimthompson5910 (jim_thompson5910):

when going from f(-2)= 3(-2) + k to 10 = 3(-2) + k I replaced f(-2) with 10 because it is given that f(-2) = 10

OpenStudy (anonymous):

ohhhhh !! okay i got it. lol thank you again

jimthompson5910 (jim_thompson5910):

sure thing